(lim xtends to zero) 1-cos(2*x)*sqrt(cos(4*x))/x^2

Grade:12

3 Answers

pvn s
11 years ago

aaah, Use Lopital rule( differentiate numerator and denom twice)

Before that, simplify using ( cos2@-1)/2=COS^2@

rahul agrwl
11 years ago

ans is 2......just write  1 - cos 2x as 2 sin ^2 x ....and go ahead as limit x > 0 sin x / x =0

mohit yadav
10 years ago

1